For a standard normal distribution, determine the probabilities in parts a through d below Click here to view page 1 of the standard normal probability table Click here to view page 2 of the standard normal probability table a, Find PCs 1 55) Plz s 155)- Round to four decimal places as needed.) b. Find P(zS-1.24) Pzs-124)- Round to four decimal places as needed) e. Find P(-083szs 1.73) P-083s2s173) FIR (Round to four decimal places as needed ) d. Find P(o 37 szs2 15) P/O 37 szs2 15)- (Round to four decimal places as neoded ) -2 Enter your answor in each of the answer boxes javascriptdoExercise(2) O Type here to search 2 345
Answer #1

a) P(Z < 1.55) = 0.9394

b) P(Z < -1.24) = 0.1075

c) P(-0.83 < Z < 1.73) = P(Z < 1.73) - P(Z < -0.83) = 0.9582 - 0.2033 = 0.7549

d) P(0.37 < Z < 2.15) = P(Z < 2.15) - P(Z < 0.37) = 0.9842 - 0.6443 = 0.3399